# Angular control noise in Advanced Virgo and implications for the Einstein Telescope

### 23 Jan 2024 v1

@@ Appendix D Noise projections: the methodology
This relationship resembles the one defined in Eq. 1, where the outcome in fact represents a change in length. While $\theta_{\textnormal{TM}}$ encompasses the entire spectrum of NI movement, $d_{\textnormal{TM}}$ has been approximated as a scalar constant, which we call static offset; the same has been done for all other TMs. As explained in [3], while an accurate representation of this process involves the convolution of the BSM spectrum with the angular motion spectrum, the predominance of low-frequency components in the BSM allows us to treat the misalignment as if it were static. The static offset value can be calculated as the RMS of the BSM. The way in which the detector responds to the length variations defined in $z$ is defined by $H\in\mathcal{R}^{4}$, which represents the detector response to a differential change in the lengths of the arm cavities. The dot product between the vectors $H$ and $z$ is the signal produced in the output channel of the detector $h$, which has been calculated using Finesse.
The amount of angular control noise calculated through the model is shown in Fig. 8 in terms of detector sensitivity. In terms of the methodology used to generate these noise projections with Finesse, the process comprises two fundamental steps. First, external disturbances are propagated through the model to calculate the motion of the test masses and subsequently determine the beam spot motion, as shown in Fig. 7. The RMS of the BSM is then used as the static offset, which in our case has been determined to be $d_{\textnormal{RMS}}=0.15$ mm. Next, this beam offset is applied to all the TMs, and the external disturbances are propagated again through the model to calculate the differential change in the length of the detector’s arms. In particular, during this second step, only sensor noise has been propagated because it is the dominant noise in the detection frequency band ($f>10$ Hz), which is now our focus; see Fig. 13. Each trace in Fig. 8 refers to a specific degree of freedom. This projection was obtained by propagating the sensor noise used to control that particular DoF. For example, the ${}_{+}$ curve was obtained by propagating the ${}_{+}$ QPD sensing noise. The last point to consider is how the static offset has been allocated to each TM. The proposed noise projection methodology takes into account the coherence between the movements of all mirrors. For example, when the ${}_{+}$ QPD sensing noise is propagated, the TMs rotate simultaneously in phase. This results in a noise projection depending on the arrangement of the static offset on each TM. For example, if the beam spot is offset in such a way that the beams in both cavities are shifted, the in-phase motion of the TMs does not change the cavity length, and therefore, does not lead to the generation of angular noise. The static offset has been distributed among the TMs to maximize the coupling of each DoF to the detector output. Each curve in Fig. 8 corresponds to a different distribution of the direction of the static offset. For example, in the case of the ${}_{+}$ curve, the matrix $D$ was defined as:

### 04 Mar 2024 v2

@@ Appendix D Noise projections: the methodology
This relationship resembles the one defined in Eq. 1, where the outcome, in fact, represents a change in length. While $\theta_{\textnormal{TM}}$ encompasses the entire spectrum of NI movement, $d_{\textnormal{TM}}$ has been approximated as a scalar constant, which we call static offset; the same has been done for all other TMs. As explained in [4], while an accurate representation of this process involves the convolution of the BSM spectrum with the angular motion spectrum, the predominance of low-frequency components in the BSM allows us to treat the misalignment as if it were static. The static offset value can be calculated as the RMS of the BSM. The way in which the detector responds to the length variations defined in $z$ is defined by $H\in\mathcal{R}^{4}$, which represents the detector response to a differential change in the lengths of the arm cavities. The dot product between the vectors $H$ and $z$ is the signal produced in the detector output channel $h$, which has been calculated using Finesse.
@@
The amount of angular control noise calculated through the AdV model is shown in Fig. 1. In terms of the methodology used to generate these noise projections, as well as for all the noise projections presented in this work, the process comprises two fundamental steps. First, external disturbances are propagated through the model to calculate the motion of the test masses and subsequently determine the beam spot motion, as shown in Fig. 20. The RMS of the BSM is then used as the static offset, which in our case has been determined to be $d_{\textnormal{RMS}}=0.15$ mm. Next, this beam offset is applied to all the TMs, and the external disturbances are propagated again through the model to calculate the differential change in the length of the detector’s arms. In particular, during this second step, only sensor noise has been propagated because it is the dominant noise in the detection frequency band ($f>10$ Hz), which is now our focus; see Fig. 12. Fig. 1 shows the total level of angular noise produced by all the DoFs under control. The contribution of each of the DoF is shown in Fig. 21. Each contribution was obtained by propagating the sensor noise used to control that particular DoF. For example, the ${}_{+}$ curve was obtained by propagating the ${}_{+}$ QPD sensing noise. The last point to consider is how the static offset has been allocated to each TM. The used noise projection methodology takes into account the coherence between the movements of all mirrors. For example, when the ${}_{+}$ QPD sensing noise is propagated, the TMs rotate simultaneously in phase. This results in a noise projection depending on the arrangement of the static offset on each TM. For example, if the beam spot is offset in such a way that the beams in both cavities are shifted, the in-phase motion of the TMs does not change the cavity length and, therefore, does not lead to the generation of angular noise. The static offset has been distributed among the TMs to maximize the coupling of each DoF to the detector output. Each curve in Fig. 21 corresponds to a different distribution of the direction of the static offset. For example, in the case of the ${}_{+}$ curve, the matrix $D$ was defined as:
Figure 21: The figure displays the angular control noise generated by each DoF under control. The methodology used to produce this projection is explained in Appendix D. The arrangement of the static offsets on each of the TMs is shown in Tab. 4. This arrangement ensures that the coupling between the controlled DoF and the detector output is maximized. In black, the target sensitivity for AdV during O3 is shown.